A 41 g bullet moving a horizontal velocity of 380 m/s comes to a stop 14 cm within a solid wall. (a) What is the change in the bullet's mechanical energy? (b) What is the magnitude of the average force from the wall stopping it?
Answer #1

a)

here

delta E = delta K = - 0.5 * m * v0^2

delta E =- 0.5 * 0.041 * 380^2

delta E = 2.96 * 10^3 J

b)

then force = delta E / d

F = 2.96 * 10^3 / 0.14

F = 2.1 * 10^4 N